# DGInStyle: Domain-Generalizable Semantic Segmentation with Image Diffusion Models and Stylized Semantic Control

### 05 Dec 2023 v1

@@ Abstract
Large, pretrained latent diffusion models (LDMs) have demonstrated an extraordinary ability to generate creative content, specialize to user data through few-shot fine-tuning, and condition their output on other modalities, such as semantic maps. However, are they usable as large-scale data generators, e.g., to improve tasks in the perception stack, like semantic segmentation? We investigate this question in the context of autonomous driving, and answer it with a resounding ”yes”. We propose an efficient data generation pipeline termed DGInStyle. First, we examine the problem of specializing a pretrained LDM to semantically-controlled generation within a narrow domain. Second, we design a Multi-resolution Latent Fusion technique to overcome the bias of LDMs towards dominant objects. Third, we propose a Style Swap technique to endow the rich generative prior with the learned semantic control. Using DGInStyle, we generate a diverse dataset of street scenes, train a domain-agnostic semantic segmentation model on it, and evaluate the model on multiple popular autonomous driving datasets. Our approach consistently increases the performance of several domain generalization methods, in some cases by +2.5 mIoU compared to the previous state-of-the-art method without our generative augmentation scheme. Source code and dataset are available at dginstyle.github.io.
Figure 1: Crossing domain boundaries with DGInStyle. We propose a data-centric generative pipeline for domain generalization. It is derived from Stable Diffusion and augmented with a novel high-precision style-preserving semantic control. DGInStyle combines semantic masks (Query) with style prompts (e.g., ”day”, ”night”) to generate training data for semantic segmentation networks with widely varying appearance. DGInStyle achieves state-of-the-art semantic segmentation across domains in autonomous driving.
@@ 1 Introduction
The rise of generative image modeling has been a game changer for AI-assisted creativity. Moreover, it also paves the way for improvements beyond artistic generation, particularly in computer vision. In this paper, we investigate one such avenue and use a powerful text-to-image generative diffusion model to improve semantic segmentation.
Segmenting images into semantically defined categories requires large annotated datasets of images and associated label maps, as a basis for supervised training. Manual annotation to obtain those label maps is time-consuming and expensive [11, 55], which is where image generation comes into play. Synthetic datasets are annotated by construction and therefore cheap to collect, but they invariably suffer from a domain gap [16], meaning that a network trained on such data (the source domain) will perform poorly on the real images of interest (the target domain). When the characteristics of the target domain are known in advance through (labeled or unlabeled) samples, the domain gap can be addressed with Domain Adaptation techniques [16, 26]. A more challenging, arguably equally important setting is Domain Generalization (DG) [76, 28, 14], where a model is deployed in a new environment without the chance to first collect data and adapt. I.e., the target domain is unknown except for the high-level application context (such as “autonomous driving”).
In the literature concerned with domain shift, the role of the prior domain is sometimes overlooked. In end-to-end pipelines, that prior typically remains implicit; for instance, it could stem from pretrained backbone weights used in most segmentation DG methods (often from ImageNet [12]) or loss functions that depend on feature space distances [26, 76]. We hypothesize that the rich prior that emerges in modern foundational models trained on internet-scale datasets [57] can further boost domain generalization. To verify this intuition, we set out to design DGInStyle, a data generation pipeline with a pretrained text-to-image LDM [50] at its core, fine-tuned with data from the source domain, with conditioning on the associated label maps. Such a pipeline can automatically generate images with characteristics of the prior domain and equipped with pixel-aligned label maps (Fig. 1). The hope is that a model trained on such data will offer improved domain generalization, drawing on the prior knowledge embedded in the LDM; which would validate our hypothesis.
Several modifications are necessary to turn an off-the-shelf LDM [50] into a data generation pipeline: First we introduce a Style Swap technique inspired by modern fine-tuning and semantic style control mechanisms, to achieve the necessary level of control and diversity over the outputs. Second, we present a Multi-Resolution Latent Fusion technique, which helps us to go beyond the limited resolution of the LDM generator, and at the same time makes it possible to appropriately handle small instances, a crucial capability for downstream performance. Third, we use the resulting data generator to train off-the-shelf semantic segmentation networks, including methods built specifically to mitigate the impact of domain shifts, and boost the state-of-the-art for domain generalization in autonomous driving by +2.5 percent mIoU (Fig. 2).
@@ 2 Related Work
Deep neural networks require extensive training data, which can be costly and time-consuming to acquire. Data access and usage scenarios are severely regulated in some domains, such as medical imaging. To mitigate this, there has been a growing interest in synthetic datasets. Due to the inevitable domain gap between synthetic datasets and application scopes, domain adaptation methods focusing on a single target, or domain generalization, focusing on the wider task-specific domain, come to the rescue.
Creating a realistic synthetic dataset often involves physically-based simulators (e.g., renderers [49]), which also turns out expensive, and a challenge in its own right. Thus, the recent trend of leveraging generative models for realistic data generation is winning in cost efficiency.
@@ 2.1 Generative Models
Diffusion Models (DMs) [25, 13, 59, 39, 58, 7] have recently demonstrated state-of-the-art image generation quality, which happened thanks to the simplified training objective, enabling training on large-scale datasets. These models involve a forward diffusion process that adds noise into data distribution and a learned reverse process that recovers the original data. To reduce the computational requirements, latent diffusion models (LDMs) [50] operate in the latent space, thus enabling absorption of internet-scale data [57]. Additionally, advances in image captioning and language conditioning, such as CLIP [46], enabled text-guided control of the generation process. These advancements suggest the emergence of strong scene-understanding prior, which can be utilized for in-domain data generation.
Despite their large size, fine-tuning LDMs is possible with a minimal computational footprint, as shown in DreamBooth [52]. To further enhance the controllability beyond text prompts, a variety of diffusion models [24, 40, 72, 30, 75, 17, 20] integrate additional inputs of condition signals to provide more granular control. Ke et al. [32] repurpose an LDM for monocular depth estimation through fine-tuning while conditioning on an image. For instance, using segmentation masks to guide generation has been a focal point of research, with methodologies primarily falling into condition-based [72, 40, 30] and guidance-based categories [2, 70, 67, 9]. The limited resolution of LDMs can be an obstacle for large-scale high-resolution data generation, yet it can also be worked around, as studied in panorama generation literature [3]. These techniques offer precise pixel-wise control and serve as a valuable means of generating image-label pairs for downstream tasks.
@@ 2.2 Dataset Generation
Approaches to paired image-mask dataset generation can be categorized into three groups. The first approach falls into the category of grounded generation [37, 35, 66, 62, 18], which generates masks with the help of a separate segmentation decoder. This often involves a pretrained off-the-shelf network, and the domain it is trained on introduces additional biases bleeding into the overall generation process. The second approach falls under the umbrella of image-to-image translation techniques. Peng et al. [45] use a DM to progressively transform images from the synthetic source domain into images resembling the target domain, guided by the source domain masks. The third approach uses semantic masks to guide the image generation (Semantic guidance[68, 72, 40]. While arguably cleaner, it also comes with caveats: the unknown distribution of masks and the degree of agreement between the generation result and the mask condition. DGInStyle follows into the last category. We use masks from the source domain and enforce the generation fidelity using the proposed Multi-Resolution Latent Fusion technique.
@@ 2.3 Domain Generalization
Domain Generalization (DG) aims to enhance the robustness of models trained on source domains and enable them to perform well on unseen domains belonging to the same task group. Compared to UDA, no data from the target domain is available along with the domain itself; it is vaguely defined through several task-specific proxy evaluation datasets. To improve domain generalization in semantic segmentation, prior methods utilize transformations such as instance normalization [42] or whitening [10] to align various source domain data with a standardized feature space. Another line of research [76, 71, 43, 77, 31, 33] focuses on domain randomization, which augments the source domain with diverse styles. For instance, Zhao et al. [76] selects basis styles from the source distribution, enabling the model to generate diverse samples during training. HGFormer [14] improves the robustness of networks by introducing a hierarchical grouping mechanism that groups pixels to form part-level and whole-level masks for class label prediction. Fig. 2 shows the progress in domain generalization over recent years measured on the task of autonomous driving scene segmentation. Improvements achieved with our approach and SOTA techniques are clearly demonstrated.
@@ 2.4 Diffusion Models for Domain Generalization
Beyond the aforementioned approaches, recent works have explored the use of diffusion models for domain generalization in semantic segmentation. Gong et al. [18] investigate how well diffusion-pretrained representations generalize to new domains and introduce a prompt randomization strategy to enhance cross-domain performance. DatasetDM [62] presents a generic dataset generation model capable of producing diverse images and corresponding perception annotations using diffusion models. Both methods implement grounded generation by training a segmentation decoder to achieve image-mask alignment. They also rely on text guidance to a large extent. Our approach takes a different semantic guidance route, exhibiting higher controllability.
@@ 3 Methods
Domain generalization for semantic segmentation aims to learn a model that is robust to unseen task domains using only annotated source domain data. In this work, given the labeled source domain data represented as $\mathbf{D}^{S}=\left\{\left(x_{i}^{S},y_{i}^{S}\right)\right\}_{i=1}^{N_{s}}$, the goal is to generalize the semantic segmentation model $f_{\theta}$ to the unseen target domain data $\mathbf{D}^{T}$, by utilizing the generated labeled dataset $\mathbf{D}^{G}=\left\{\left(x_{i}^{G},y_{i}^{G}\right)\right\}_{i=1}^{N_{g}}$. In these notations, $x$ and $y$ stand for the images and their corresponding labels, respectively, whereas $N_{s}$ and $N_{g}$ are the total number of images in each dataset. The set $\{y_{i}^{G}\}_{i=1}^{N_{g}}$ is a subset of $\{y_{i}^{S}\}_{i=1}^{N_{s}}$ in our case, although other labels are possible.
@@ 3.1 Label Conditioned Image Generation
Figure 6: Style variations. DGInStyle can generate various scene conditions using style prompting. The same condition mask is used to illustrate the pixel-perfect alignment of the generations.
@@ 3.2 Preserving Style Prior with Style Swap
DreamBooth [52] was originally proposed as an efficient protocol for few-shot fine-tuning of a pretrained LDM to learn a new concept, represented in the training images and unavailable in the prior domain. We propose to use it to fine-tune the LDM towards whole domains.
@@ 4 Experiments
Our model is based on Stable Diffusion 1.5 [50] and requires a single consumer-grade GPU for training.
@@ 4.2 Implementation Details
DG Method
DGInStyle
CS [11] BDD [69] MV [41] Avg3 ACDC [55] DZ [54] Avg5
[…]
@@
+1.2 +3.1 +1.3 +1.9 +2.0 +4.6 +2.5
Table 1: DG with GTA and MiT-B5 [65]. Comparison of Domain Generalization (DG) methods for semantic segmentation in autonomous driving scenes w/ and w/o integrating our generated dataset (mIoU $\uparrow$ in %) with GTA as source domain and MiT-B5 backbone. As seen, leveraging our proposed data generation pipeline, which exploits rich generative priors and semantic conditioning, provides a substantial boost in performance across various configurations.
@@ 4.3 Comparison with State-of-the-Art DG
[…]
@@
DG Method
DGInStyle
CS [11] BDD [69] MV [41] Avg3 ACDC [55] DZ [54] Avg5
[…]
IBN-Net [42] 37.37 34.21 36.81 36.13 - - -
DRPC [71] 42.53 38.72 38.05 39.77 - - -
ISW [10] 37.20 33.36 35.57 35.38 - - -
FSDR [29] 44.80 41.20 43.40 43.13 24.77 9.66 32.77
GTR [43] 43.70 39.60 39.10 40.80 - - -
SAN-SAW [44] 45.33 41.18 40.77 42.23 - - -
AdvStyle [77] 44.51 39.27 43.48 42.42 - - -
SHADE [76] 46.66 43.66 45.50 45.27 29.06 8.01 34.58
HRDA [27] 39.63 38.69 42.21 39.18 26.08 7.80 30.28
HRDA [27] 46.89 42.81 50.19 46.63 34.19 16.16 38.05
Table 2: DG with GTA and ResNet-101 [21]. Comparison of Domain Generalization (DG) methods w/ and w/o integrating our generated dataset (mIoU $\uparrow$ in %) with GTA source domain and ResNet-101 backbone.
Furthermore, we extend our comparison to a range of DG methods using a ResNet-101 [21] backbone. We adapt HRDA for this setting by substituting their original backbones with ResNet-101. As shown in Tab. 2, our generated dataset enhances performance on the majority of the target datasets, establishing new state-of-the-art results. This indicates that our dataset generation pipeline is effective across different architectures.
To broaden the scope of our evaluation, we also leverage Cityscapes [11] as the source domain, generalizing to other real-world domains. Models are trained on Cityscapes and tested on BDD, MV, ACDC, and DZ. The results in this real-to-real adaptation scenario, presented in Tab. 3, again confirm that our generated dataset consistently boosts the performance of semantic segmentation models across all configurations. These findings affirm the versatility and robustness of our method.
DG Method
DGInStyle
BDD [69] MV [41] ACDC [55] DZ [54] Average
[…]
@@
+2.1 +1.0 2.6 +0.3 +1.5
HRDA [27, 28] 58.49 68.32 59.70 31.07 54.40
58.84 67.99 61.00 32.60 55.10
+0.4 -0.3 +1.3 +1.5 +0.7
Table 3: DG with Cityscapes and MiT-B5 [65]. Cityscapes to other datasets domain generalization w/ and w/o integrating our generated dataset (mIoU $\uparrow$ in %) with MiT-B5 backbone.

### 08 Apr 2024 v2

@@ Abstract
Large, pretrained latent diffusion models (LDMs) have demonstrated an extraordinary ability to generate creative content, specialize to user data through few-shot fine-tuning, and condition their output on other modalities, such as semantic maps. However, are they usable as large-scale data generators, e.g., to improve tasks in the perception stack, like semantic segmentation? We investigate this question in the context of autonomous driving, and answer it with a resounding ”yes”. We propose an efficient data generation pipeline termed DGInStyle. First, we examine the problem of specializing a pretrained LDM to semantically-controlled generation within a narrow domain. Second, we propose a Style Swap technique to endow the rich generative prior with the learned semantic control. Third, we design a Multi-resolution Latent Fusion technique to overcome the bias of LDMs towards dominant objects. Using DGInStyle, we generate a diverse dataset of street scenes, train a domain-agnostic semantic segmentation model on it, and evaluate the model on multiple popular autonomous driving datasets. Our approach consistently increases the performance of several domain generalization methods compared to the previous state-of-the-art methods. Source code and dataset are available at dginstyle.github.io.
Figure 1: Crossing domain boundaries with DGInStyle. We propose a data-centric generative pipeline for domain generalization. It is derived from Stable Diffusion and augmented with a novel high-precision style-preserving semantic control. DGInStyle combines semantic masks (Query) with style prompts (e.g., Night or Rain) to generate training data for semantic segmentation networks with widely varying appearance. It achieves state-of-the-art semantic segmentation across domains in autonomous driving.
@@ 1 Introduction
The rise of generative image modeling has been a game changer for AI-assisted creativity. Moreover, it also paves the way for improvements beyond artistic generation, particularly in computer vision. In this paper, we investigate one such avenue and use a powerful text-to-image generative diffusion model to improve the robustness of semantic segmentation with respect to domain shifts.
Semantic segmentation requires large annotated datasets for supervised training. While manual annotation is costly and time-consuming [11, 55], synthetic datasets offer a cost-effective solution. However, these datasets face a domain gap [16], leading to poor performance when networks trained on the source domain are applied to the target domain. When the characteristics of the target domain are known, the domain gap can be addressed with Domain Adaptation techniques [16, 26]. A more challenging, arguably equally important setting is Domain Generalization (DG) [76, 28, 14], where a model is deployed in a new environment without knowing the target domain except for its general context (such as “autonomous driving”).
In the DG semantic segmentation literature, the role of the prior domain is often overlooked. In end-to-end pipelines, that prior typically remains implicit; for instance, it could stem from pretrained backbone weights used in most segmentation DG methods (often from ImageNet [12]) or loss functions that depend on feature space distances [28, 76]. Therefore, we take a closer look at the prior domain and study how we can utilize the rich prior that emerges in modern foundational models trained on internet-scale datasets [57] to improve domain generalization of semantic segmentation. To this end, we design DGInStyle, a novel data generation pipeline with a pretrained foundational text-to-image LDM [50] at its core, fine-tuned with data from the source domain, with conditioning on the associated dense label maps. Such a pipeline can automatically generate images with characteristics of the prior domain and equipped with pixel-aligned label maps (Fig. 1). Armed with such a pipeline, we approach DG differently from other methods by focusing on synthesizing data instead of model architectures or training techniques. The idea is that a model trained on such data will offer improved domain generalization, drawing on the prior knowledge embedded in the LDM.
@@
This comes with two new challenges: The LDM needs to learn to produce images that match semantic masks from the labeled source domain while avoiding overfitting to its style. Additionally, the generated images need precise alignment with segmentation masks, even for very small instances. Our Contributions address these issues: First we propose a novel Style Swap technique inspired by modern fine-tuning and semantic style control mechanisms, to achieve the necessary level of control and diversity over the outputs. It is based on the novel finding that the semantic control and the underlying (stylized) diffusion model can be decoupled and swapped. This enables our simple yet efficient Style Swap, which allows to learn dense semantic control on the source domain while removing the undesired source domain style. Second, we present a novel Multi-Resolution Latent Fusion technique, which helps us to go beyond the limited resolution of the LDM generator. It is an essential step to achieve conditioned generation of small instances, which is crucial for learning semantic segmentation on generated data. Without it, a segmentation model trained on it would struggle with inconsistent generated images and segmentation masks. Lastly, we use the resulting generative pipeline to create a diversified dataset to train semantic segmentation networks, including methods to mitigate the impact of domain shifts. Due to its complementary design, DGInStyle achieves major performance improvements when combined with existing DG methods. In particular, it significantly boosts the state-of-the-art domain generalization in autonomous driving, as shown in Fig. 2.
@@ 2 Related Work
Deep neural networks require extensive training data, which can be costly and time-consuming to acquire. Data access and usage scenarios are severely regulated in some domains, such as medical imaging. To mitigate this, there has been a growing interest in synthetic datasets. Due to the inevitable domain gap between synthetic datasets and application scopes, domain adaptation methods focusing on a single target domain, or domain generalization, focusing on the wider task-specific domain, come to the rescue. Creating a realistic synthetic dataset often involves physically-based simulators (e.g., renderers [49]), which also turns out expensive, and a challenge in its own right. Thus, the recent trend of leveraging generative models for realistic data generation is winning in cost efficiency.
@@ 2.1 Generative Models
Diffusion Models (DMs) [25, 13, 59, 39, 58, 7] have recently demonstrated state-of-the-art image generation quality, which happened thanks to the simplified training objective, enabling training on large-scale datasets. These models involve a forward diffusion process that adds noise into data and a learned reverse process that recovers the original data. To reduce the computational requirements, latent diffusion models (LDMs) [50] operate in the latent space, thus enabling absorption of internet-scale data [57]. Additionally, advances in image captioning and language conditioning, such as CLIP [46], enabled text-guided control of the generation process. These advancements suggest the emergence of strong scene-understanding prior, which can be utilized for in-domain data generation.
Despite their large size, DreamBooth [52] demonstrated that LDMs can be efficiently fine-tuned. To further enhance the controllability beyond text prompts, a variety of diffusion models [24, 40, 72, 30, 75, 17, 20] integrate additional inputs of condition signals to provide more granular control. As demonstrated in [32], LDMs can be repurposed to learn new tasks through fine-tuning and extra conditioning. The usage of segmentation masks to guide generation has been a focal point of research, with methodologies primarily falling into condition-based [72, 40, 30] and guidance-based categories [2, 70, 67, 9]. When using pretrained off-the-shelf models, the limited resolution of LDMs can be an obstacle to large-scale high-resolution data generation. Yet, it can also be worked around, as studied in panorama generation literature [3]. These techniques offer precise pixel-wise control and, subsequently, a means of generating image-label pairs for downstream tasks.
@@ 2.2 Dataset Generation
Approaches to paired image-mask dataset generation can be categorized into three groups. The first approach falls into the category of grounded generation [37, 35, 66, 62, 18], which generates masks with the help of a separate segmentation decoder. This often involves a pretrained off-the-shelf network, and the domain it is trained on introduces additional biases bleeding into the overall generation process. The second approach falls under the umbrella of image-to-image translation techniques. [45] use a DM to progressively transform images from the synthetic source domain into images resembling the target domain, guided by the source domain masks. The third approach uses semantic masks to guide the image generation (Semantic guidance[68, 72, 40]. While arguably cleaner, it also comes with caveats: the unknown distribution of masks and the degree of agreement between the generation result and the mask condition. DGInStyle follows into the last category. We use masks from the source domain and enforce the generation fidelity using the proposed Multi-Resolution Latent Fusion technique.
@@ 2.3 Domain Generalization
Domain Generalization (DG) aims to enhance the robustness of models trained on source domains and enable them to perform well on unseen domains belonging to the same task group. Compared to UDA, no data from the target domain is available along with the domain itself; it is defined through a union of task-specific proxy evaluation datasets. To improve domain generalization in semantic segmentation, prior methods utilize transformations such as instance normalization [42] or whitening [10] to align various source domain data with a standardized feature space. Another line of research [76, 71, 43, 77, 31, 33] focuses on domain randomization, which augments the source domain with diverse styles. For instance, [76] selects basis styles from the source distribution, enabling the model to generate diverse samples during training. HGFormer [14] improves the robustness of networks by introducing a hierarchical grouping mechanism that groups pixels to form part-level and whole-level masks for class label prediction. Fig. 2 shows the progress in domain generalization over recent years measured on the task of autonomous driving scene segmentation. Improvements achieved with our approach and SOTA techniques are clearly demonstrated.
@@ 2.4 Diffusion Models for Domain Generalization
Beyond the aforementioned approaches, recent works have explored the use of diffusion models for domain generalization in semantic segmentation. Gong et al. [18] investigate how well diffusion-pretrained representations generalize to new domains and introduce a prompt randomization strategy to enhance cross-domain performance. DatasetDM [62] presents a generic dataset generation model capable of producing diverse images and corresponding annotations using diffusion models. These methods implement grounded generation by training a segmentation decoder to achieve image-mask alignment. Our approach takes a different semantic guidance route, exhibiting higher controllability and generating consistent image-label pairs that qualify as training datasets.
@@ 3 Methods
Domain generalization for semantic segmentation aims to learn a model that is robust to unseen task domains using only annotated source domain data. In this work, given the labeled source domain represented as $\mathbf{D}^{\mathcal{S}}=\left\{\left(x_{i}^{\mathcal{S}},y_{i}^{\mathcal{S}}\right)\right\}_{i=1}^{N_{\mathcal{S}}}$, the goal is to generalize the semantic segmentation model $f_{\theta}$ to unseen target domains $\mathbf{D}^{\mathcal{T}}$ from the same task group, by utilizing the generated labeled dataset $\mathbf{D}^{\mathcal{G}}=\left\{\left(x_{i}^{\mathcal{G}},y_{i}^{\mathcal{G}}\right)\right\}_{i=1}^{N_{\mathcal{G}}}$ in style of the prior domain $\mathcal{P}$ (hence DGInStyle), thus maximizing the overlap with the target domain. In these notations, $x$ and $y$ stand for the images and their corresponding labels, respectively, whereas $N_{\mathcal{S}}$ and $N_{\mathcal{G}}$ are the total number of images in each dataset. The set $\{y_{i}^{\mathcal{G}}\}_{i=1}^{N_{\mathcal{G}}}$ is a subset of $\{y_{i}^{\mathcal{S}}\}_{i=1}^{N_{\mathcal{S}}}$ in our case, although other labels are possible.
@@ 3.1 Label Conditioned Image Generation
(a) Car, Rider…
(b) Foggy
(c) Rainy
(d) Snowy
Figure 6: Style variations. DGInStyle can generate images under various scene conditions through style prompting, while maintaining consistent dense semantic control from (a).
@@ 3.2 Preserving Style Prior with Style Swap
DreamBooth [52] was originally proposed as an efficient protocol for few-shot fine-tuning of a pretrained LDM to learn a new concept, represented in the training images and unavailable in the prior domain. We employ its reconstruction loss as an efficient means for fine-tuning the LDM towards whole domains.
@@ 4.2 Implementation Details
DG Method DGInStyle CS [11] BDD [69] MV [41] Avg3 ACDC [55] DZ [54] Avg5 $\Delta$Avg5
ResNet-101 [21]
IBN-Net [42] 37.37 34.21 36.81 36.13 25.85 6.12 28.07 $\uparrow$ 5.1
40.80 38.98 43.20 40.99 31.68 11.19 33.17
RobustNet [10] 37.20 33.36 35.57 35.38 24.80 5.49 27.28 $\uparrow$ 6.8
41.03 39.62 44.85 41.83 32.30 12.73 34.11
DRPC [71] 42.53 38.72 38.05 39.77
FSDR [29] 44.80 41.20 43.40 43.13 24.77 9.66 32.77
GTR [43] 43.70 39.60 39.10 40.80
SAN-SAW [44] 45.33 41.18 40.77 42.23
AdvStyle [77] 44.51 39.27 43.48 42.42
SHADE [76] 46.66 43.66 45.50 45.27 29.06 8.01 34.58
HRDA [27, 28] 39.63 38.69 42.21 40.18 26.08 7.80 30.88 $\uparrow$ 7.2
46.89 42.81 50.19 46.63 34.19 16.16 38.05
MiT-B5 [65]
Color-Aug 46.64 45.45 49.04 47.04 36.10 16.37 38.72 $\uparrow$ 3.3
@@
Table 1: DG with GTA source domain and ResNet-101/MiT-B5 backbone. Comparison of Domain Generalization (DG) methods for semantic segmentation in autonomous driving scenes w/ and w/o integrating our generated dataset (mIoU $\uparrow$ in %) with GTA as source domain, using either ResNet-101 or MiT-B5 as the backbone. As seen, leveraging our proposed data generation pipeline, which exploits rich generative priors and semantic conditioning, provides a substantial boost in performance across various configurations.
@@ 4.3 Comparison with State-of-the-Art DG
In Tab. 1, we benchmark several DG methods trained using either the GTA dataset alone or augmented with our DGInStyle and subsequently evaluated across five real-world datasets to measure their generalization from GTA to other domains. Specifically, we integrate DGInStyle into IBN-Net [42], RobustNet [10], Color-Aug (random brightness, contrast, saturation, and hue), DAFormer [26, 28], and HRDA [27, 28] covering CNN-based ResNet-101 [21] and Transformer-based MiT-B5 [65] network architectures.
The results in Tab. 1 indicate that DGInStyle significantly enhances the DG performance across various DG methods and network architectures. The improvements range from +2.5 mIoU up to +7.2 mIoU on the average over 5 datasets. In particular, DGInStyle improves the overall state-of-the-art performance by a significant gain of +2.5 mIoU. These results confirms the efficacy of our method in generating diverse, style-varied image-label pairs for semantic segmentation, thereby significantly contributing to robust domain generalization across different network architectures and training strategies.
@@
To broaden the scope of our evaluation, we set an experiment with Cityscapes [11] as a source domain, generalizing to other real-world domains in Tab. 2. As a real-world dataset, Cityscapes has a smaller domain gap to the other real-world target datasets than the synthetic GTA dataset. When using Cityscapes as a source, the baseline performance without DGInStyle is therefore naturally higher, which reduces the potential for improvement. Yet, even in this more saturated setting, DGInStyle achieves significant average improvements. These findings affirm the versatility and robustness of our method.
DG Method DGInStyle BDD [69] MV [41] ACDC [55] DZ [54] Average $\Delta$Average
Color-Aug 53.33 60.06 52.38 23.00 47.19 $\uparrow$ 2.1
@@
HRDA [27, 28] 58.49 68.32 59.70 31.07 54.40 $\uparrow$ 0.7
58.84 67.99 61.00 32.60 55.11
Table 2: DG with Cityscapes source domain and MiT-B5 [65] backbone. Cityscapes to other datasets domain generalization w/ and w/o integrating our generated dataset (mIoU $\uparrow$ in %).